OronSH:

A hunter and an invisible rabbit play a game in the Euclidean plane. The rabbit's starting point, $A_0,$ and the hunter's starting point, $B_0$ are the same. After $n-1$ rounds of the game, the rabbit is at point $A_{n-1}$ and the hunter is at point $B_{n-1}.$ In the $n^{\text{th}}$ round of the game, three things occur in order: i. The rabbit moves invisibly to a point $A_n$ such that the distance between $A_{n-1}$ and $A_n$ is exactly $1.$ ii. A tracking device reports a point $P_n$ to the hunter. The only guarantee provided by the tracking device to the hunter is that the distance between $P_n$ and $A_n$ is at most $1.$ iii. The hunter moves visibly to a point $B_n$ such that the distance between $B_{n-1}$ and $B_n$ is exactly $1.$ Is it always possible, no matter how the rabbit moves, and no matter what points are reported by the tracking device, for the hunter to choose her moves so that after $10^9$ rounds, she can ensure that the distance between her and the rabbit is at most $100?$

ChooseAUsename:

trivial

OronSH:

but can you prove it

ChooseAUsename:

Yes

OronSH:

what is your proof

ChooseAUsename:

bump again

OronSH:

where is your proof \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] bruh imagien dollar signs don't work for latex

OronSH:

how do i type inline latex bruh

OronSH:

bruh \$\frac12+\frac13=\frac56\$ why does this not work

ChooseAUsename:

Consider a positive integer \(n>d\), to be chosen later. Let the hunter start at \(B\) and the rabbit at \(A\), as shown. Let \(\ell\) denote line \(AB\). Now, we may assume the rabbit reveals its location \(A\), so that all previous information becomes irrelevant. The rabbit chooses two points \(X\) and \(Y\) symmetric about \(l\) such that \(XY=2\) and \(AX=AY=n\), as shown. The rabbit can then hop to either \(X\) or \(Y\), pinging the point \(P_n\) on the \(\ell\) each time. This takes \(n\) hops. Now among all points \(H\) the hunter can go to, \(\min \max\{HX,HY\}\) is clearly minimized with \(H\in\ell\) by symmetry. So the hunter moves to a point \(H\) such that \(BH=n\) as well. In that case the new distance is \(HX=HY\). We now compute \begin{align*}HX^2&=1+HM^2=1+(\sqrt{AX^2+1}-AH)^2\\&=1+(\sqrt{n^2-1}-(n-d))^2\\&\ge1+\left(\left(n-\frac1n\right)-(n-d)\right)^2\\&=1+(d-1/n)^2\end{align*} which exceeds \(d^2+\frac12\) whenever \(n\ge4d\).\(\qquad\square\)

OronSH:

oh wait does \(this work\) o it's \(

OronSH:

@chooseausename wrote:
Consider a positive integer \(n>d\), to be chosen later. Let the hunter start at \(B\) and the rabbit at \(A\), as shown. Let \(\ell\) denote line \(AB\). Now, we may assume the rabbit reveals its location \(A\), so that all previous information becomes irrelevant. The rabbit chooses two points \(X\) and \(Y\) symmetric about \(l\) such that \(XY=2\) and \(AX=AY=n\), as shown. The rabbit can then hop to either \(X\) or \(Y\), pinging the point \(P_n\) on the \(\ell\) each time. This takes \(n\) hops. Now among all points \(H\) the hunter can go to, \(\min \max\{HX,HY\}\) is clearly minimized with \(H\in\ell\) by symmetry. So the hunter moves to a point \(H\) such that \(BH=n\) as well. In that case the new distance is \(HX=HY\). We now compute \[\begin{align*}HX^2&=1+HM^2=1+(\sqrt{AX^2+1}-AH)^2\\&=1+(\sqrt{n^2-1}-(n-d))^2\\&\ge1+\left(\left(n-\frac1n\right)-(n-d)\right)^2\\&=1+(d-1/n)^2\end{align*}\] which exceeds \(d^2+\frac12\) whenever \(n\ge4d\).\(\qquad\square\)
bruh sobad you didn't even copy evan chen's solution properly

OronSH:

you didn't even finish the problem

Fortish:

@oronsh wrote:
you didn't even finish the problem
He can guide you to the answer but he can't give it directly.

Fortish:

@oronsh wrote:
oh wait does \(this work\) o it's \(
please remain on topic

OronSH:

ok but where's evan's diagram

ChooseAUsename:

Extrinix:

At least source your inputs, please. Evan Chen - IMO Notes

OronSH:

bruh not even asymptote so bad

OronSH:

@extrinix wrote:
At least source your inputs, please. Evan Chen - IMO Notes
ok

Extrinix:

@oronsh wrote:
@extrinix wrote:
At least source your inputs, please. Evan Chen - IMO Notes
ok
Talking to @chooseausename

OronSH:

ok

OronSH:

@chooseausename wrote:
ok

ChooseAUsename:

@extrinix wrote:
@oronsh wrote:
@extrinix wrote:
At least source your inputs, please. Evan Chen - IMO Notes
ok
Talking to @chooseausename
ok

ChooseAUsename:

Fortish:

Please reread the rules or if you didn't read the rules: https://questioncove.com/rules

Fortish:

@thisgirlpretty are you able to lock this question, ppl keep spamming.

Fortish:

@vocaloid can you lock.